# Timescales diagnostics for saving viscous and MHD-driven dusty discs from external photoevaporation

### v1

@@ Abstract
Methods. We run a suite of 1D simulations of smooth discs with varying initial sizes, different levels of viscous and MHD-wind stresses modeled via an $\alpha$ parametrisation, and strengths of the external FUV environment. We then track disc properties such as their radii, various lifetime diagnostics, and the amount of dust removed by the photoevaporative wind, as a function of the underlying parameters.
@@ 1 Introduction
Despite its importance, external photoevaporation of protoplanetary discs is a poorly constrained process. The count of star forming regions (SFRs) with few stars (say, $\lesssim$ few hundred) is high, so these SFRs are commonly found close to Earth, are easily observed, and have been thoroughly covered by surveys (e.g. Andrews et al. 2013, 2018, Ansdell et al. 2016, 2018, Tazzari et al. 2021, Pascucci et al. 2016, Barenfeld et al. 2016, Long et al. 2018, Cieza et al. 2019, Garufi et al. 2021, Zhang et al. 2025). These regions have initial mass functions (IMFs) skewed towards low-mass K and M dwarfs where more massive stars (O/B) are less frequent. This makes them especially suitable for high-resolution campaigns (e.g. with the Atacama Large Millimeter Array or ALMA), since their proximity ($\lesssim 300\,\mathrm{pc}$) and relative isolation minimizes contamination and extinction, and requires shorter observation times. On the other hand, they paint a rather biased picture of the environments in which protoplanetary discs evolve. As they host few massive stars, they are not subject to high external FUV fluxes, which are thus limited to $\sim 10\,G_{0}$ (Winter et al. 2018, Anania et al. 2025a, 2025b; here, $G_{0}$ is the Habing unit commonly used to measure FUV fluxes, Habing 1968). However, SFRs with high number of stars (e.g., Orion, Carina, or Cygnus OB2) host within them such a large number of them that a given star taken at random across all SFRs is more likely to have been born with a massive star nearby (e.g. Fatuzzo and Adams 2008). In other words, the typical protoplanetary disc observed in surveys so far is not representative of the planet-hosting star in the galaxy in terms of its neighbouring environment, where the FUV flux may exceed $100-1000\,G_{0}$.
@@
On the theoretical side, previous works have modeled the response of the gas to FUV fields (Clarke 2007), including its dependence on the internal stresses (Coleman et al. 2024), others have investigated what happens to dust under external irradiation (e.g., Sellek et al. 2020), or how theoretical predictions for purely viscous dusty discs within mild FUV environments relate to observations (Anania et al. 2025a). However, no study has systematically described the evolution of both the gas and the dust in viscously- vs. MHD-driven discs that are subject to external irradiation. This is the main aim of this paper. The specific question we tackle is under which conditions the dust component can survive long enough so that there is enough time for planet formation to occur. To answer this question, we perform 1D simulations of the dust and gas components of discs with various internal stresses and different external FUV fluxes, and extract disc lifetimes that relate to both the gas and the dust.
@@
The rest of the paper is organised as follows. In Section 2 we describe the disc model and the parameter space investigated in this work. In Section 3 we present the main results for the evolution of gaseous and solid components. In Section 4 we discuss their relevance in the context of our current understanding of disc evolution, disc observations, and the assembly of planetary systems, and we discuss the limitations of our modeling. Finally, we conclude in Section 5.
@@ 2.1 Surface density evolution and angular momentum transport
We now turn to the parameters $\alpha_{\mathrm{SS}}$ and $\alpha_{\mathrm{DW}}$ that appear in eq. (1). Following the approach of Shakura and Sunyaev (1973), these quantities conveniently parametrize the (time-averaged) radial stress tensor (describing the radial transport of angular momentum) and the vertical stress tensor (describing the extraction of angular momentum by the MHD wind) respectively, including both Reynolds and Maxwell stresses [^1] These $\alpha$ parameters effectively measure stress-to-pressure ratios, with the explicit definitions (Tabone et al. 2022b) $\alpha_{\mathrm{SS}}:=\frac{2}{3}\frac{\int_{-H_{\mathrm{DW}}}^{+H_{\mathrm{DW}}}T_{r\phi}\,\mathrm{d}z}{\Sigma_{\mathrm{g}}c_{\mathrm{s}}^{2}},\quad\alpha_{\mathrm{DW}}:=\frac{4}{3}\frac{r\left[T_{z\phi}\right]_{-H_{\mathrm{DW}}}^{+H_{\mathrm{DW}}}}{\Sigma_{\mathrm{g}}c_{\mathrm{s}}^{2}},$ where $T_{r\phi}$ and $T_{z\phi}$ are the components of the (time-averaged) stress tensor associated to the radial and vertical transport of angular momentum, and $H_{\mathrm{DW}}$ is the vertical scale height where the MHD wind is launched. . Following Tabone et al. (2022b), we assume constant $\alpha_{\mathrm{SS}}$ and $\alpha_{\mathrm{DW}}$ for simplicity throughout. We then define a total $\alpha$ parameter
@@ 2.2 External photoevaporation
Our treatment of external photoevaporation follows closely that of Sellek et al. (2020) and Anania et al. (2025a). We prescribe a photoevaporative mass-loss rate $\dot{\Sigma}_{\mathrm{g},\mathrm{evap}}(r)$ in eq. (1) as follows. We use the updated FRIEDv2 grid of Haworth et al. (2023) with a fiducial ratio between the Polycyclic Aromatic Hydrocarbon (PAH) abundance and the dust abundance of 1 (see Appendix A for a quantitative comparison with the original FRIED grid from Haworth et al. 2018, used e.g. in Sellek et al. 2020). Given a fixed $M_{*}$ and an FUV flux, we obtain a local mass-loss rate $\dot{\tilde{M}}_{\mathrm{g},\mathrm{evap},i}$ at each radius $r_{i}$ (where $r_{i}$ are the radial cell centers used in the code as described in Subsect. 2.4), using a bilinear interpolation of the FRIEDv2 grid over $r_{i}$ and $\Sigma_{\mathrm{g}}(r_{i})$. This $\dot{\tilde{M}}_{\mathrm{g},\mathrm{evap},i}$ is, however, not the actual local mass-loss rate that should be used. This is because, as shown by Sellek et al. (2020), while the photoevaporation rate first increases with fixed $r$ in the weakly-bound regions that are optically thick in the FUV and then decrease with $r$ in the optically thin region, the flow in the outermost, optically thin disc region region is actually set by the larger rates from the closer-in, optically thick disc regions. Thus, simply applying the local mass-loss rate from the FRIEDv2 grid taken at face value would be unphysical because the local mass-loss rate in the optically thin region assumes that the region interior to it does not flow in the wind, which is in reality not the case. Sellek et al. (2020) showed that the actual mass-loss rate is set at (and slightly outside of) the radius located at the transition between the optically thick and optically thin regimes. This so-called truncation radius $R_{\mathrm{t}}$ (cfr. also Anania et al. 2025a) is the radial location which maximises the mass-loss rate $\dot{\tilde{M}}_{\mathrm{g},\mathrm{evap},i}$ (the one obtained from the FRIED grid assuming the radial cell centre $r_{i}$ to be the outer radius of the disc, Sellek et al. 2020). Mass removal by external FUV can thus be implemented inside-out, at radii $r>R_{\mathrm{t}}$ (i.e., over the radial grid, at radii $r_{i}$ with $i>i_{\mathrm{t}}$), obtaining first a total mass-loss rate given by
$\dot{M}_{\mathrm{g},\mathrm{tot},\mathrm{evap}}=\sum_{i>i_{\mathrm{t}}}\dot{\tilde{M}}_{\mathrm{g},\mathrm{evap},i}\frac{M_{\mathrm{g},i}}{M_{\mathrm{g}}(r>R_{\mathrm{t}})},$ (8)
and then spreading this locally at larger radii in proportion to the local mass. The effective local mass-loss rate is thus
@@ 2.3 Dust component
assuming spherical particles with radius $a_{k}$ and material density $\rho_{\mathrm{s}}$, which we take to be $1\,\mathrm{g}\,\mathrm{cm}^{-3}$. The diffusion coefficient $D_{\mathrm{d},k}=\frac{D_{\mathrm{g}}}{\mathrm{Sc}_{k}}$, where the Schmidt number $\mathrm{Sc}_{k}=1+\mathrm{St}_{k}^{2}$ (Youdin and Lithwick 2007), and we take $D_{\mathrm{g}}=\nu_{\mathrm{g}}=\nu_{\mathrm{SS}}$, simply the gas turbulent viscosity. Dust feedback onto the gas is included as in Booth et al. (2017) and 2020MNRAS.492.1279S, although it makes little difference due to small dust mass fractions.
@@ 3.1 Gas evolution
For the gas component, the coupling of purely viscously-evolving discs (VE discs for short) with external photoevaporation was modelled in Clarke (2007), who showed that these discs follow a predictable evolution. As long as the accretion rate onto the star is larger than the mass-loss rate at the photoevaporative edge, the disc expands outwards and self-adjusts to a state where the outward viscous flow at the outer edge matches $\dot{M}_{\mathrm{evap}}$. When $\dot{M}_{\mathrm{acc}}$ is no longer larger than the photoevaporative rate, the disc outer edge moves inward on a timescale comparable to the local viscous time-scale. Thus, VE discs experience a long-lived phase where the accretion rate onto the star matches the photoevaporative rate; they also lose memory of the initial disc cutoff radius $R_{\mathrm{c}}$} (Sellek et al. 2020). The evolution is expected to be different in the MHD-wind case (DW discs for short): these discs do not expand outwards (e.g. review by Manara et al. 2023), there is no restoring mechanism to reach a self-adjustment between the inner and outer regions, and the MHD-disc-wind mass removal mechanism is decoupled from external photoevaporation. Because there is no natural equilibrium state, a pure DW disc is then not independent of initial radius $R_{\mathrm{c}}$} at later times. These trends are shown in Figure 1 for a highly irradiated disc, $1000\,G_{0}$, and with moderate total $\alpha_{\mathrm{tot}}=10^{-3}$ (top row), as a function of different initial disc radii.
@@
Another consequence of this is that, for a purely DW disc, the long-lived phase of disc evolution has $\dot{M}_{\mathrm{acc}}>\dot{M}_{\mathrm{evap}}$. This is because either $\dot{M}_{\mathrm{acc}}>\dot{M}_{\mathrm{evap}}$ from the start (for compact discs), or (for larger disc) $\dot{M}_{\mathrm{evap}}$ quickly drops within less than 1 Myr, as material is not replenished at larger radial separations. Therefore, observing discs at random will preferentially reveal them at this evolutionary stage. Instead, sampling VE discs will typically reveal a $\dot{M}_{\mathrm{acc}}-\dot{M}_{\mathrm{evap}}$ correlation. This explains the recent results from Weder et al. (2025), which propose distinguishing between the two angular momentum transport mechanisms by looking at accretion and evaporative mass loss rates (they specifically focused on Cygnus OB2, with FUV fluxes of order $10^{2}-10^{4}$; Fig. 1 uses fluxes of $10^{3}\,G_{0}$ for comparison). In particular, in the viscous case we expect an almost 1:1 relation between the two rates, while this is not the case for the (pure) MHD wind model. We note, however, that this distinction is almost completely lost in the hybrid model ($\psi_{\mathrm{DW}}=1$), where the little viscosity $\alpha_{\mathrm{SS}}$ is sufficient to recover the viscous spreading effects. To drive this point further, we perform mock-population synthesis experiments where we sample discs with a fixed $\psi_{\mathrm{DW}}$, across (finite) FUV fluxes, $\alpha_{\mathrm{tot}}$, and $R_{\mathrm{c}}$} randomly chosen across their respective ranges, performing “observations” at random times with a uniform distribution $\mathcal{U}_{[0\,\mathrm{My},31.6\,\mathrm{My}]}$. We plot in Figure 2 the observed $\dot{M}_{\mathrm{acc}}$ vs. $\dot{M}_{\mathrm{wind}}$, defined as either $\dot{M}_{\mathrm{wind}}=\dot{M}_{\mathrm{evap}}$ or $\dot{M}_{\mathrm{wind}}=\dot{M}_{\mathrm{evap}}+\dot{M}_{\mathrm{DW}}$. Although very simplistic, this experiment shows how the $\dot{M}_{\mathrm{acc}}-\dot{M}_{\mathrm{evap}}$ correlation disappears with higher and higher $\psi_{\mathrm{DW}}$.
@@ 3.2 Dust evolution
Dust evolution, in both the VE and DW cases, is driven by growth, radial drift and removal from external photoevaporation, and these processes compete with one another in making dust lifetimes longer or shorter. Dust that is well-coupled to the gas can follow its outward expansion in the viscous case. Similarly, dust diffusion is active when $\psi_{\mathrm{DW}}\ll 1$ but not in the MHD-wind case (because $D_{\mathrm{d}}=\nu_{\mathrm{SS}}/{\mathrm{Sc}}$, cfr. subsect. 2.3). Since the photoevaporative mass-loss rate increases with radius in the optically-thick regime (Sellek et al. 2020), external irradiation would cause DW discs to show longer lifetimes than their purely viscous counterparts. Instead, drift causes dust to move inward and accrete onto the star. When $\psi_{\mathrm{DW}}\gg 1$, because dust does not diffuse outward (like it does in the viscous case) it also has less of a radial displacement to cover; the relative efficiency of drift and gas drag will depend on the Stokes number. The fact that DW discs do not diffuse outwards also means that their surface density profiles in the outer disc will be steeper as they are carved out by external FUV irradiation, so that inward drift can be more efficient the outer regions depending on the disc’s initial size. Just like for the gas, an important difference is in the discs’ ability to expand via diffusion, which is fundamentally different for VE and DW discs (see Zagaria et al. 2022). But, just like the gas (Coleman et al. 2024), this difference can be strongly suppressed by high external FUV fluxes, which can obfuscate the distinction between the different angular momentum transport mechanisms.

### v2

@@ Abstract
Methods. We run a suite of 1D simulations of smooth discs with varying initial sizes, different levels of viscous and MHD-wind stresses modelled via an $\alpha$ parametrisation, and strengths of the external FUV environment. We then track disc properties such as their radii, various lifetime diagnostics, and the amount of dust removed by the photoevaporative wind, as a function of the underlying parameters.
@@ 1 Introduction
Despite its importance, external photoevaporation of protoplanetary discs is a poorly constrained process. The count of star forming regions (SFRs) with few stars (say, $\lesssim$ few hundred) is high, so these SFRs are commonly found close to Earth, are easily observed, and have been thoroughly covered by surveys (e.g. Andrews et al. 2013, 2018, Ansdell et al. 2016, 2018, Tazzari et al. 2021, Pascucci et al. 2016, Barenfeld et al. 2016, Long et al. 2018, Cieza et al. 2019, Garufi et al. 2021, Zhang et al. 2025). These regions have initial mass functions (IMFs) skewed towards low-mass K and M dwarfs where more massive stars (O/B) are less frequent. This makes them especially suitable for high-resolution campaigns (e.g. with the Atacama Large Millimeter Array or ALMA), since their proximity ($\lesssim 300\,\mathrm{pc}$) and relative isolation minimises contamination and extinction, and requires shorter observation times. On the other hand, they paint a rather biased picture of the environments in which protoplanetary discs evolve. As they host few massive stars, they are not subject to high external FUV fluxes, which are thus limited to $\sim 10\,G_{0}$ (Winter et al. 2018, Anania et al. 2025a, 2025b; here, $G_{0}$ is the Habing unit commonly used to measure FUV fluxes, Habing 1968). However, SFRs with high number of stars (e.g., Orion, Carina, or Cygnus OB2) host within them such a large number of them that a given star taken at random across all SFRs is more likely to have been born with a massive star nearby (e.g. Fatuzzo and Adams 2008). In other words, the typical protoplanetary disc observed in surveys so far is not representative of the planet-hosting star in the galaxy in terms of its neighbouring environment, where the FUV flux may exceed $100-1000\,G_{0}$.
@@
On the theoretical side, previous works have modelled the response of the gas to FUV fields (Clarke 2007), including its dependence on the internal stresses (Coleman et al. 2024), others have investigated what happens to dust under external irradiation (e.g., Sellek et al. 2020), or how theoretical predictions for purely viscous dusty discs within mild FUV environments relate to observations (Anania et al. 2025a). However, no study has systematically described the evolution of both the gas and the dust in viscously- vs. MHD-driven discs that are subject to external irradiation. This is the main aim of this paper. The specific question we tackle is under which conditions the dust component can survive long enough so that there is enough time for planet formation to occur. To answer this question, we perform 1D simulations of the dust and gas components of discs with various internal stresses and different external FUV fluxes, and extract disc lifetimes that relate to both the gas and the dust.
@@
The rest of the paper is organised as follows. In Section 2 we describe the disc model and the parameter space investigated in this work. In Section 3 we present the main results for the evolution of gaseous and solid components. In Section 4 we discuss their relevance in the context of our current understanding of disc evolution, disc observations, and the assembly of planetary systems, and we discuss the limitations of our modelling. Finally, we conclude in Section 5.
@@ 2.1 Surface density evolution and angular momentum transport
We now turn to the parameters $\alpha_{\mathrm{SS}}$ and $\alpha_{\mathrm{DW}}$ that appear in eq. (1). Following the approach of Shakura and Sunyaev (1973), these quantities conveniently parametrise the (time-averaged) radial stress tensor (describing the radial transport of angular momentum) and the vertical stress tensor (describing the extraction of angular momentum by the MHD wind) respectively, including both Reynolds and Maxwell stresses [^1] These $\alpha$ parameters effectively measure stress-to-pressure ratios, with the explicit definitions (Tabone et al. 2022b) $\alpha_{\mathrm{SS}}:=\frac{2}{3}\frac{\int_{-H_{\mathrm{DW}}}^{+H_{\mathrm{DW}}}T_{r\phi}\,\mathrm{d}z}{\Sigma_{\mathrm{g}}c_{\mathrm{s}}^{2}},\quad\alpha_{\mathrm{DW}}:=\frac{4}{3}\frac{r\left[T_{z\phi}\right]_{-H_{\mathrm{DW}}}^{+H_{\mathrm{DW}}}}{\Sigma_{\mathrm{g}}c_{\mathrm{s}}^{2}},$ where $T_{r\phi}$ and $T_{z\phi}$ are the components of the (time-averaged) stress tensor associated to the radial and vertical transport of angular momentum, and $H_{\mathrm{DW}}$ is the vertical scale height where the MHD wind is launched. . Following Tabone et al. (2022b), we assume constant $\alpha_{\mathrm{SS}}$ and $\alpha_{\mathrm{DW}}$ for simplicity throughout. We then define a total $\alpha$ parameter
@@ 2.2 External photoevaporation
Our treatment of external photoevaporation follows closely that of Sellek et al. (2020) and Anania et al. (2025a). We prescribe a photoevaporative mass-loss rate $\dot{\Sigma}_{\mathrm{g},\mathrm{evap}}(r)$ in eq. (1) as follows. We use the updated FRIEDv2 grid of Haworth et al. (2023) with a fiducial ratio between the Polycyclic Aromatic Hydrocarbon (PAH) abundance and the dust abundance of 1 (see Appendix A for a quantitative comparison with the original FRIED grid from Haworth et al. 2018, used e.g. in Sellek et al. 2020). Given a fixed $M_{*}$ and an FUV flux, we obtain a local mass-loss rate $\dot{\tilde{M}}_{\mathrm{g},\mathrm{evap},i}$ at each radius $r_{i}$ (where $r_{i}$ are the radial cell centres used in the code as described in Subsect. 2.4), using a bilinear interpolation of the FRIEDv2 grid over $r_{i}$ and $\Sigma_{\mathrm{g}}(r_{i})$. This $\dot{\tilde{M}}_{\mathrm{g},\mathrm{evap},i}$ is, however, not the actual local mass-loss rate that should be used. This is because, as shown by Sellek et al. (2020), the photoevaporation rate first increases with $r_{i}$ in the weakly-bound regions that are optically thick in the FUV and then decrease with $r_{i}$ in the optically thin region. However, the flow in the outermost, optically thin disc region is actually set by the larger rates from the closer-in, optically thick disc regions. Thus, simply applying the local mass-loss rate from the FRIEDv2 grid at $r_{i}$, taken at face value, would be unphysical. This is because the local mass-loss rate in the optically thin region assumes that the region interior to it does not flow in the wind, which is in reality not the case. Sellek et al. (2020) showed that the actual mass-loss rate is set at (and slightly outside of) the radius located at the transition between the optically thick and optically thin regimes. This so-called truncation radius $R_{\mathrm{t}}$ (cfr. also Anania et al. 2025a) is the radial location which maximises the mass-loss rate $\dot{\tilde{M}}_{\mathrm{g},\mathrm{evap},i}$ (the one obtained from the FRIED grid assuming the radial cell centre $r_{i}$ to be the outer radius of the disc, Sellek et al. 2020). Mass removal by external FUV can thus be implemented inside-out, at radii $r>R_{\mathrm{t}}$ (i.e., over the radial grid, at radii $r_{i}$ with $i>i_{\mathrm{t}}$). First, we obtain a total mass-loss rate given by
$\dot{M}_{\mathrm{g},\mathrm{tot},\mathrm{evap}}=\sum_{i>i_{\mathrm{t}}}\dot{\tilde{M}}_{\mathrm{g},\mathrm{evap},i}\frac{M_{\mathrm{g},i}}{M_{\mathrm{g}}(r>R_{\mathrm{t}})}.$ (8)
Then, this mass-loss rate is distributed at larger radii in proportion to the local mass. The effective local mass-loss rate is thus
@@ 2.3 Dust component
assuming spherical particles with radius $a_{k}$ and material density $\rho_{\mathrm{s}}$, which we take to be $1\,\mathrm{g}\,\mathrm{cm}^{-3}$. The diffusion coefficient $D_{\mathrm{d},k}=\frac{D_{\mathrm{g}}}{\mathrm{Sc}_{k}}$, where the Schmidt number $\mathrm{Sc}_{k}=1+\mathrm{St}_{k}^{2}$ (Youdin and Lithwick 2007), and we take $D_{\mathrm{g}}=\nu_{\mathrm{g}}=\nu_{\mathrm{SS}}$, simply the gas turbulent viscosity. Dust feedback onto the gas is included as in Booth et al. (2017) and Sellek et al. (2020), although it makes little difference due to small dust mass fractions.
@@ 3.1 Gas evolution
For the gas component alone, the coupling of purely viscously-evolving discs (VE discs for short) with external photoevaporation has been modelled for the first time by Clarke (2007), who showed that these discs follow a predictable evolution. As long as the accretion rate onto the star is larger than the mass-loss rate at the photoevaporative edge, the disc expands outwards and self-adjusts to a state where the outward viscous flow at the outer edge matches $\dot{M}_{\mathrm{evap}}$. When $\dot{M}_{\mathrm{acc}}$ is no longer larger than the photoevaporative rate, the disc outer edge moves inward on a timescale comparable to the local viscous time-scale. Thus, VE discs experience a long-lived phase where the accretion rate onto the star matches the photoevaporative rate; they also lose memory of the initial disc cutoff radius $R_{\mathrm{c}}$} (Sellek et al. 2020). The evolution is expected to be different in the MHD-wind case (DW discs for short): these discs do not expand outwards (e.g. review by Manara et al. 2023), there is no restoring mechanism to reach a self-adjustment between the inner and outer regions, and the MHD-disc-wind mass removal mechanism is decoupled from external photoevaporation. Because there is no natural equilibrium state, a pure DW disc is then not independent of initial radius $R_{\mathrm{c}}$} at later times. These trends are shown in Figure 1 for a highly irradiated disc, $1000\,G_{0}$, and with moderate total $\alpha_{\mathrm{tot}}=10^{-3}$ (top row), as a function of different initial disc radii.
@@
Another consequence of the qualitatively different behaviour in VE vs. DW discs is that, for a purely DW disc, the long-lived phase of disc evolution has $\dot{M}_{\mathrm{acc}}>\dot{M}_{\mathrm{evap}}$. This is because either $\dot{M}_{\mathrm{acc}}>\dot{M}_{\mathrm{evap}}$ from the start (for compact discs), or (for larger disc) $\dot{M}_{\mathrm{evap}}$ quickly drops within less than 1 Myr, as material is not replenished at larger radial separations. Therefore, observing discs at random will preferentially reveal them at this evolutionary stage. Instead, sampling VE discs will typically reveal a $\dot{M}_{\mathrm{acc}}-\dot{M}_{\mathrm{evap}}$ correlation. This explains the recent results from Weder et al. (2025), which propose distinguishing between the two angular momentum transport mechanisms by looking at accretion and evaporative mass loss rates (they specifically focused on Cygnus OB2, with FUV fluxes of order $10^{2}-10^{4}$; Fig. 1 uses fluxes of $10^{3}\,G_{0}$ for comparison). In particular, in the viscous case we expect an almost 1:1 relation between the two rates, while this is not the case for the (pure) MHD wind model. We note, however, that this distinction is almost completely lost in the hybrid model ($\psi_{\mathrm{DW}}=1$), where the little viscosity $\alpha_{\mathrm{SS}}$ is sufficient to recover the viscous spreading effects. To drive this point further, we perform mock-population synthesis experiments where we sample discs with a fixed $\psi_{\mathrm{DW}}$, across (finite) FUV fluxes, $\alpha_{\mathrm{tot}}$, and $R_{\mathrm{c}}$} randomly chosen across their respective ranges, performing “observations” at random times with a uniform distribution $\mathcal{U}_{[0\,\mathrm{My},31.6\,\mathrm{My}]}$. We plot in Figure 2 the observed $\dot{M}_{\mathrm{acc}}$ vs. $\dot{M}_{\mathrm{wind}}$, defined as either $\dot{M}_{\mathrm{wind}}=\dot{M}_{\mathrm{evap}}$ or $\dot{M}_{\mathrm{wind}}=\dot{M}_{\mathrm{evap}}+\dot{M}_{\mathrm{DW}}$. Although very simplistic, this experiment shows how the $\dot{M}_{\mathrm{acc}}-\dot{M}_{\mathrm{evap}}$ correlation disappears with higher and higher $\psi_{\mathrm{DW}}$.
@@ 3.2 Dust evolution
Dust evolution, in both the VE and DW cases, is driven by growth, radial drift and removal from external photoevaporation, and these processes compete with one another in making dust lifetimes longer or shorter. Dust that is well-coupled to the gas can follow its outward expansion in the viscous case. Similarly, dust diffusion is active when $\psi_{\mathrm{DW}}\ll 1$ (VE discs) but not in the MHD-wind case (because $D_{\mathrm{d}}=\nu_{\mathrm{SS}}/{\mathrm{Sc}}$, cfr. subsect. 2.3). Since the photoevaporative mass-loss rate increases with radius in the optically-thick regime (Sellek et al. 2020), external irradiation would cause DW discs to show longer lifetimes than their purely viscous counterparts. Instead, drift causes dust to move inward and accrete onto the star. When $\psi_{\mathrm{DW}}\gg 1$ (DW discs), because dust does not diffuse outward (like it does in the viscous case) it also has less of a radial displacement to cover; the relative efficiency of drift and gas drag will depend on the Stokes number. The fact that DW discs do not diffuse outwards also means that their surface density profiles in the outer disc will be steeper as they are carved out by external FUV irradiation, so that inward drift can be more efficient the outer regions depending on the disc’s initial size. Just like for the gas, an important difference is in the discs’ ability to expand via diffusion, which is fundamentally different for VE and DW discs (see Zagaria et al. 2022). But, just like the gas (Coleman et al. 2024), this difference can be strongly suppressed by high external FUV fluxes, which can obfuscate the distinction between the different angular momentum transport mechanisms.